OpenStudy (anonymous):

what is the total number of integers between 100 and 200 that are divisible by 3? steps needed

OpenStudy (anonymous):

102 is the first, now count in 3s, 102,105,108.... the largest is 198. so there are (198-102)/3+1=32+1=33

OpenStudy (owlfred):

ah that is very helpful

OpenStudy (anonymous):

thank you :)

OpenStudy (owlfred):

rsvitale, I just gave you a medal for your help!

OpenStudy (anonymous):

thanks owlfred/ you're welcome dhiwagm :)